# Ab initio study of Proximity-Induced Superconductivity in PbTe/Pb heterostructures

### v1

@@ 4 Conclusion
We used a combined Kohn–Sham and Bogoliubov–de Gennes approach to investigate the semiconductor–superconductor (SM–SC) heterostructure composed of PbTe and Pb, focusing on the interfacial properties and the potential applicability of this system for detecting Majorana Zero Modes (MZMs) 19, 9, 7. We discussed the (normal state) metallization of the heterostructure driven by strong hybridization between PbTe and Pb states at the interface. We predict a significant difference in electrostatic potential at the interface, which poses a challenge in achieving the necessary energy alignment between the SM Zeeman gap and SC gap. These findings remain robust against strain and electric field.
@@
In the superconducting state, a soft BCS–like superconducting gap is computed, indicating a weak superconducting coupling between the two systems 10. Moreover, we predicted proximity–induced superconductivity in the PbTe region of the heterostructure. We resolved the superconducting density of states layer by layer, revealing that the induced SC gap is enhanced near the interface, with a softened SC gap on the Pb side. We analyzed the shape and decay of the anomalous charge density $\chi(\mathbf{r})$, revealing an anisotropic pairing potential. We observed a weak superconducting coupling regime at the interface, which prevents the superconductor to overscreen the HS properties. The weak regime favours the device’s tunability with gate and bias voltages, creating a route to get the desirable energy alignment between the two materials. Our findings strengthen the understanding of the physics governing PbTe/Pb hybrid devices and their potential application in hosting Majorana Zero Modes (MZMs).
@@ Acknowledgments
The authors acknowledge the fruitful discussion with Erik Bakkers, Arnold H. Kole and Nils Wittemeier. ZZ acknowledges the research program “Materials for the Quantum Age” (QuMat) for financial support. This program (registration number 024.005.006) is part of the Gravitation program financed by the Dutch Ministry of Education, Culture and Science (OCW). RR and ZZ acknowledge financial support from Sector Plan Program 2019-2023. This work was sponsored by NWO-Domain Science for the use of supercomputer facilities. We also acknowledge that the results of this research have been achieved using the Tier-0 PRACE Research Infrastructure resource Discoverer based in Sofia, Bulgaria (OptoSpin project id. 2020225411). This project has received funding from the European Union’s Horizon Europe research and innovation program under Grant Agreement No 101130384 (QUONDENSATE).
Supporting Information for:
Ab initio study of Proximity-Induced Superconductivity in PbTe/Pb heterostructures
R. Reho, Andrés R. Botello-Méndez, and Zeila Zanolli
The PDF file includes:
Normal and superconducting properties of bulk Pb and PbTe, theoretical methods, details concerning the initialization of the superconducting pairing potential, geometrical aspects of the heterostructures, and supplementary figures.
S5 Supplementary note 1: Computational methods
To model the PbTe/Pb interface, we computed its electronic structure using density functional theory (DFT) as implemented in the SIESTA code 20, 14, including spin-orbit coupling and the PBE exchange-correlation functional 21. We employed optimized norm-conserving Vanderbilt pseudopotentials (ONCVPSP 22) in PSML format 23 from PseudoDojo database 24. The Kohn-Sham equations were solved using standard double-zeta polarized basis sets of localized atomic orbitals on a $60\times 60\times 1$ Monkhorst-Pack grid. Real space integral were evaluated on grids with a cut-off energy of 2000 Ry for the SC state calculations and 500  Ry for the normal state ones. The Fermi-Dirac occupation function was smoothed using an electronic temperature of 0.1 meV, to ensure that the broadening of the electronic states is well below the size of the superconducting gap. All crystal structures were relaxed using the FIRE algorithm with a force threshold of 0.005 eV/Å and a maximum stress tolerance of 0.5 meV/Å3.
The superconducting properties of PbTe/Pb were modeled using the SIESTA–BdG method 13. In this approach a semi-empirical superconducting pairing potential is introduced on top of the self-consistent Kohn-Sham Hamiltonian describing the normal state. We used the fixed–$\Delta$ method with the superconducting pairing potential initialized in real space as a touching spherical hardwells with radius $\mathbf{r}=1.58$ and strength $\bar{\Delta}=1.50$ meV around Pb atoms. We disregard the singlet–triplet mixing effect 25 as the computed contributions from triplet states are considerably lower in magnitude than the singlet ones.
Figure S6: Orbital (a) and spin (b) projected bands structure of bulk Pb in the normal state. In (a) we observe that the electronic states crossing the Fermi level are $p$ and $d$ orbitals. In (b) we observe a spin flip at $W$ for $S_{x}$ and $S_{y}$ spin orientations.
S6 Supplementary note 2: Geometrical aspect of the HS
@@
We constructed the PbTe/Pb HS, stacking 15 layers slab of PbTe cell and 11 layers of Pb along the [001] direction. The HS extends infinitely in the x and y directions. Periodic replicae are separated by 26 Å  vacuum along the z–axis. The PbTe slab was constructed by rotating the in-plane conventional cell by 45 degrees and reconstructing its primitive cell. The Pb slab was constructed by repeating the conventional cell along the [001] direction 12 times and removing one layer to achieve 11 layers of Pb. Farey sequences 26 were employed to identify commensurate cells for the full heterostructure. This method inherently introduces strain in the system. We considered several strained configurations of the PbTe/Pb heterostructure. The strain on each side of the heterostructure is defined as:
$\text{strain}=\frac{\mathrm{a}_{\mathrm{hs}}-\mathrm{a}_{\mathrm{rlx,i}}}{\mathrm{a}_{\mathrm{rlx,i}}}\quad\mathrm{i=Pb,PbTe}$ (S1)
@@
S7 Supplementary note 3: Bulk Pb and bulk PbTe
@@
Figure S7: (a) SC–DOS (continous lines) and ADOS (dashed lines) of bulk Pb computed with the solution methods: non SCF-BdG (black), Fixed-$\Delta$ (blue), and Full SCF-BdG (red). The Full SCF-BdG method reveals a small splitting of the main coherence peaks. (b) Orbital contribution to the DOS using the Full SCF-BdG method, showing that the main contributions from $p$ and $d$ orbitals, as indicated by the orbitals crossing the Fermi level. Similar results are found using the other two solution methods.
S7.1 Bulk Pb
@@
We modeled the superconducting state using the SIESTA-BdG code 13, employing three solution methods: non SCF-BdG, Fixed-$\Delta$, and full SCF-BdG. All three solution methods predict a conventional U-shaped SC gap of $\sim$3 meV (Figure S7). We initialize the pairing potential in the superconducting strength representation (on a real-space grid) as touching spherical hardwells with radius $\mathbf{r}=1.58$. In the non SCF–BdG and fixed–$\Delta$ methods, the initial strength of the superconducting pairing potential is $\bar{\Delta}=1.50$ meV. In the full SCF-BdG method, instead, the initial value of the superconducting coupling is $\lambda(\mathbf{r})=80.3$ eV. We computed the orbital projected superconducting density of states (Figure S7) and conclude that Pb $p$ and $d$ are the dominant electronic states close to the Fermi level, and hence relevant for the superconducting properties.
S7.2 Bulk PbTe
@@
Table 2: Electronic gap ($\Delta E_{L}$) of PbTe as a function of its in-plane lattice parameter. The configuration that matches the value of the lattice parameter for the (4,-5) PbTe/Pb HS is highlighted in bold.
@@
Figure S8: Electronic band structures for PbTe bulk. (a) Orbital projected band structure. The inset shows an energy region close to the Fermi level. (b) Spin projected band structure. Positive and negative spins are depicted in red and blue, respectively.
@@
Figure S9: Projected DOS of bulk PbTe. Near the Fermi level, the valence band is predominantly influenced by Te $p$ orbitals, whereas the conduction band shows a substantial contribution from Pb $p$ orbitals
@@
S8 Supplementary note 4: Initialization of the Pairing Potential and Its Impact on the Superconducting Properties
In the SIESTA-BdG method one can employ different methods for solving the BdG equations and different approaches for initializing the pairing potential $\Delta$ 13. We tested multiple initializations of the pairing potential. In the main text, we report results obtained within the superconducting strength representation for all simulations.
We compared the superconducting strength and orbital representation (intra–orbital coupling or full–overlap) for the PbTe/Pb heterostructure (Figure S10). The intra–orbital coupling lacks the interactions between different orbitals that can be captured by the superconducting strength representation, making it more localized. As a result, the intra–orbital coupling misses the hybridization and coupling between states belonging to the PbTe and Pb sides of the heterostructure, leading to a U–shaped DOS. However, when we excessively couple states belonging to different orbitals and atoms (full–overlap), the superconducting gap closes. Therefore, initializing the pairing potential in real–space as touching sphere hardwells is a physically intuitive and reasonable approach.
@@
Figure S10: The SC–DOS, $\rho(\varepsilon))$, of the PbTe/Pb heterostructure is computed using the superconducting strength representation method (solid black line) and the orbital representation method. In the orbital representation, we evaluate intra-orbital coupling (solid orange line) and full-overlap coupling (dashed blue line). When pairing occurs between electrons and holes within the same orbital on the same atom (intra–orbital), a conventional U-shaped gap emerges. In contrast, when superconducting coupling involves interactions between different orbitals across different atoms (superconducting strength representation or full–overlap), the hard superconducting gap becomes softened.
S9 Supplementary note 5: Mulliken charge analysis
@@
Table 3: Excess of charge $\Delta Q$ computed for different strain. Positive (negative) values of $\Delta Q$ denote the acceptor (donor) character of the layer. We compare the excess of charge on the PbTe side ($\Delta Q_{PbTe}$) of the heterostructure, with the one on the Pb side ($\Delta Q_{Pb}$). The (4, -5) HS is highlighted in bold.
@@
Figure S11: Mulliken charge analysis for PbTe/Pb. Layer-by-layer excess charge. Positive (negative) values of $\Delta Q$ denote the acceptor (donor) character of the layer. The red dashed line denote the separation between the PbTe and Pb side of the heterostructure. We notice that at the interface there is a depletion of charge on the PbTe side of the heterostructure, while the opposite is true on the Pb side. The sum of the excess charge on the PbTe and Pb side $\Delta Q$ is reported in the plot.

### v2

@@ V Conclusion
We used a combined Kohn–Sham and Bogoliubov–de Gennes approach to investigate the semiconductor–superconductor (SM–SC) heterostructure composed of PbTe and Pb, focusing on the interfacial properties. We discussed the (normal state) metallization of the heterostructure driven by strong hybridization between PbTe and Pb states at the interface. We predict a significant difference in electrostatic potential at the interface, which poses a challenge in achieving the necessary energy alignment between the SM Zeeman gap and SC gap. These findings remain robust against strain and electric field.
@@
In the superconducting state, a BCS–like superconducting gap is computed, indicating a intermediate superconducting coupling between the two systems. Moreover, we predicted proximity–induced superconductivity in the PbTe region of the HS. We resolved the superconducting density of states layer by layer, revealing that the induced SC gap is enhanced near the interface, with a poisoned SC gap on the Pb side. We analyzed the shape and decay of the anomalous charge density $\chi(\mathbf{r})$, revealing an anisotropic pairing potential. We observed a weak superconducting coupling regime at the interface, which prevents the superconductor to overscreen the HS properties. The weak regime favours the device’s tunability with gate and bias voltages, creating a route to get the desirable energy alignment between the two materials.
The field of topological quantum materials is increasingly moving towards identifying optimal design parameters to control and enhance the functionality of hybrid devices. Our results naturally fit in this direction, by providing first-principles quantitative predictions of the superconducting and electronic properties of the PbTe/Pb semiconductor/superconductor interface and their resiliance to strain and electric field. We expect our findings to benefit both experimentalists and theoreticians alike, due to the key role played by this interface in various technologically relevant devices, among which core/shell semiconductor/superconductor nanowires.
@@ Acknowledgments
The authors acknowledge the fruitful discussion with Erik Bakkers, Arnold H. Kole and Nils Wittemeier. ZZ acknowledges the research program “Materials for the Quantum Age” (QuMat) for financial support. This program (registration number 024.005.006) is part of the Gravitation program financed by the Dutch Ministry of Education, Culture and Science (OCW). RR and ZZ acknowledge financial support from Sector Plan Program 2019-2023. This work was sponsored by NWO-Domain Science for the use of supercomputer facilities. This publication is part of the project ”Quantum Materials by Design” with file number 2024.012 of the research programme Computing Time on National Computer Facilities which is (partly) financed by the Dutch Research Council (NWO) under the grant https://doi.org/10.61686/JUXDK41645 . We also acknowledge that the results of this research have been achieved using the Tier-0 PRACE Research Infrastructure resource Discoverer based in Sofia, Bulgaria (OptoSpin project id. 2020225411). This project has received funding from the European Union’s Horizon Europe research and innovation program under Grant Agreement No 101130384 (QUONDENSATE).
VI Data availability
The data of this work are available at the following NOMAD repository [30].
VII keywords
proximity interaction, superconductivity, semiconductor-superconductor heterostructures, first-principles simulations, Bogoliubov de Gennes equations
Appendix A Geometrical aspect of the HS
@@
We constructed the PbTe/Pb HS by stacking 15 layers of PbTe and 11 layers of Pb along the [001] direction. The HS extends infinitely in the x and y directions. Periodic replicae are separated by 26 Å of vacuum along the z–axis. The PbTe slab was constructed by rotating the in-plane conventional cell by 45 degrees and reconstructing its primitive cell. Farey sequences [22] were employed to identify commensurate cells for the full heterostructure. This method inherently introduces strain in the system. We considered several strained configurations of the PbTe/Pb heterostructure. The strain on each side of the heterostructure is defined as:
$\text{strain}=\frac{\mathrm{a}_{\mathrm{hs}}-\mathrm{a}_{\mathrm{rlx,i}}}{\mathrm{a}_{\mathrm{rlx,i}}}\quad\mathrm{i=Pb,PbTe}$ (A)
@@
Appendix B Bulk Pb and bulk PbTe
@@
1 Bulk Pb
@@
Figure B1: Orbital (a) and spin (b) projected band structure of bulk Pb in the normal state. In (a) we observe that the electronic states crossing the Fermi level are $p$ and $d$ orbitals. In (b) we observe a spin flip at $W$ for $S_{x}$ and $S_{y}$ spin orientations.
Figure B2: SC–DOS $\rho(\varepsilon)$ (continous lines) and ADOS $\chi(\varepsilon)$ (dashed lines) for bulk Pb computed with the solution methods (a) non SCF-BdG, (b) fixed-$\Delta$, and (c) full SCF-BdG. Panel (d) is the orbital projected normal density of states. Some lines at zero DOS are not visible as they are hidden behind by the magenta full line.
We computed the superconducting state using the SIESTA-BdG code [31], including SOC and employing three solution methods: non SCF-BdG, fixed-$\Delta$, and full SCF-BdG. All three solution methods predict a conventional U-shaped SC gap of $\sim$3 meV (Fig. B2).
In the following, we discuss the components of the normal $\rho$ and anomalous $\chi$ densities, highlighting how those are obtained from the projected densities of states $\rho^{\alpha\beta}_{\mu\nu}$ and $\chi^{\alpha\beta}_{\mu\nu}$ computed by the code.
The projected densities of states are given by [31]
$\rho_{\mu\nu}^{\alpha\beta}(\mathbf{k})=$ $\sum_{i}f\left(\varepsilon_{i}(\mathbf{k})\right)u_{i\mu}^{*\alpha}(\mathbf{k})u_{i\nu}^{\beta}(\mathbf{k})$ (B)
$+\sum_{i}\bar{f}\left(\varepsilon_{i}(-\mathbf{k})\right)v_{i\mu}^{\alpha}(-\mathbf{k})v_{i\nu}^{*\beta}(-\mathbf{k}),$
$\chi_{\mu\nu}^{\alpha\beta}(\mathbf{k})=$ $\sum_{i}f\left(\varepsilon_{i}(\mathbf{k})\right)u_{i\nu}^{\beta}(\mathbf{k})v_{n\mu}^{*\alpha}(\mathbf{k})$
$+\sum\bar{f}\left(\varepsilon_{i}(-\mathbf{k})\right)u_{i\mu}^{\alpha}(-\mathbf{k})v_{i\nu}^{*\beta}(-\mathbf{k}).$
The tensor structure describes how the weight is distributed among orbitals and spin degrees of freedom. Each projected density of states contains an ‘electron” (proportional to $f(\varepsilon_{i}({\mathbf{k}}))$) and ‘hole” (proportional to $\bar{f}(\varepsilon_{i}({\mathbf{k}}))$) contribution, corresponding to the two components of the BdG quasiparticles. The spin sector of $\rho^{\alpha\beta}_{\mu\nu}$ is expanded in the basis of Pauli matrices ($\sigma^{0},\sigma^{x},\sigma^{y},\sigma^{z}$). This yields one spin-summed term and a three components term, conventionally labelled as $x,y,z$. For the anomalous density $\chi^{\alpha\beta}_{\mu\nu}$ the same decomposition is rearranged into singlet (S) and triplet (T) components according to the Balian-Werthamer representation [5]. Summing over all orbital and spin indices one recovers the scalar density of states ($\rho(\varepsilon)$ and $\chi(\varepsilon)$).
In Fig. B2 we report the total (spin-summed) normal density as sum of the electron and hole components $\rho_{tot_{e}}+\rho_{tot_{h}}$, its $i=x,y,z$ spin-components $\rho_{i}=\rho_{e_{i}}+\rho_{h_{i}}$, and the singlet and triplet components of the anomalous density $\chi_{S}$ and $\chi_{T}$.
The z component of the SC density of states $\rho(\varepsilon)_{z}$ exhibits small oscillations around zero when computed using the non-SCF-BdG and fixed-$\Delta$ methods but vanishes when computed using the full SCF-BdG method, which is the most accurate one [Fig. B2.(a-c)]. The x and y components of the normal density are always identically zero, regardless of the solution method employed. The triplet components $\chi(\varepsilon)_{T_{x}}$ is almost zero for non SCF-BdG and fixed-$\Delta$ solution methods, and identically zero for the full SCF-BdG one. The $\chi(\varepsilon)_{T_{y}}$ and $\chi(\varepsilon)_{T_{z}}$ components of $\chi$ are always zero. The small oscillations around zero for $\rho(\varepsilon)$ and $\chi(\varepsilon)$ are due to numerical noise. The oscillations appear only for the non SCF-BdG and fixed-$\Delta$ solution methods because in these cases the anomalous components are not updated, namely they are kept constant with respect to the initial assumption. This further validates our implementation and explains why the DOS of triplet components of $\chi(\varepsilon)$ are symmetrical with respect to the $E_{F}$ in the non SCF-BdG and fixed- $\Delta$ cases.
@@
2 Bulk PbTe
[…]
Figure B3: Electronic band structures for PbTe bulk. (a) Orbital projected band structure. The inset shows an energy region close to the Fermi level. (b) Spin projected band structure. Positive and negative spins are depicted in red and blue, respectively.
Table B1: Electronic gap ($\Delta E_{L}$) of PbTe as a function of its in-plane lattice parameter. The configuration that matches the value of the lattice parameter for the (4,-5) PbTe/Pb HS is highlighted in bold.
@@ 2 Bulk PbTe
Figure B4: Projected DOS of bulk PbTe. Near the Fermi level, the valence band is predominantly influenced by Te $p$ orbitals, whereas the conduction band shows a substantial contribution from Pb $p$ orbitals
@@
Appendix C PbTe/Pb HS: Initialization of the Pairing Potential and Its Impact on the Superconducting Properties
@@
Figure C5: The SC–DOS, $\rho(\varepsilon))$, of the (4,-5) PbTe/Pb HS is computed using the superconducting strength representation method (solid black line) and the orbital representation method. In the orbital representation, we evaluate intra-orbital coupling (solid orange line) and full-overlap coupling (dashed blue line). When pairing occurs between electrons and holes within the same orbital on the same atom (intra–orbital), a conventional U-shaped gap emerges. In contrast, when superconducting coupling involves interactions between different orbitals across different atoms (superconducting strength representation or full–overlap), the hard superconducting gap becomes poisoned.
In the SIESTA-BdG method one can employ different methods for solving the BdG equations and different approaches for initializing the pairing potential $\Delta$ [31]. We tested multiple initializations of the pairing potential. In the main text, we report results obtained within the superconducting strength representation for all simulations.
We compared the superconducting strength and orbital representation (intra–orbital coupling or full–overlap) for the PbTe/Pb heterostructure (Fig. C5). The intra–orbital coupling lacks the interactions between different orbitals that can be captured by the superconducting strength representation, making it more localized. As a result, the intra–orbital coupling misses the hybridization and coupling between states belonging to the PbTe and Pb sides of the heterostructure, leading to a U–shaped DOS. However, when we excessively couple states belonging to different orbitals and atoms (full–overlap), the superconducting gap closes. Therefore, initializing the pairing potential in real–space as touching sphere hardwells is a physically intuitive and reasonable approach.
Appendix D Additional figures and tables
This sections presents additional figures and tables related to the PbTe/Pb HS. We performed a Mulliken charge analysis and calculated the excess charge $\Delta Q$ under different strain conditions, finding that $\Delta Q$ remains robust agains strain variations. (Table D1). A layer-by-layer decomposition of $\Delta Q$ (Fig. D6) reveals charge depletion on the PbTe side of the interface, with a corresponding accumulation on the Pb side. Fig. D7 shows the BdG spectrum of the (4,-5) PbTe/Pb HS. The SC gap $\Delta(\mathbf{k})$ is anisotropic, as evidenced by the variations in the width of the splitting of the BdG dispersion along the Brillouin Zone
@@ Appendix D Additional figures and tables
Figure D6: Mulliken charge analysis for (4,-5) PbTe/Pb HS. Layer-by-layer excess charge. Positive (negative) values of $\Delta Q$ denote the acceptor (donor) character of the layer. The red dashed line denote the separation between the PbTe and Pb side of the heterostructure. The sum of the excess charge on the PbTe and Pb side $\Delta Q$ is reported in the plot.
Figure D7: BdG spectrum for the (4,-5) PbTe/Pb HS. The right panel displays the superconducting states dispersion in a narrower energy range.
Figure D8: (a) SC–DOS $\rho(\varepsilon)$ and (b) ADOS $\chi(\varepsilon)$ for the (4,-5) PbTe/Pb HS computed with the fixed-$\Delta$ solution method. The triplet components of $\chi(\varepsilon)$ are clearly negligible.
For the PbTe/Pb (4,-5) HS, we compared the components of $\rho(\varepsilon)$ and $\chi(\varepsilon)$ following the approach discussed for bulk Pb in Appendix 1. Using the fixed-$\Delta$ solution method, we observed a similar behaviour as for bulk Pb and concluded that the triplet components of the anomalous density can also be neglected in the HS (Fig. D8).
Table D1: Excess of charge $\Delta Q$ computed for different strain. Positive (negative) values of $\Delta Q$ denote the acceptor (donor) character of the layer. We compare the excess of charge on the PbTe side ($\Delta Q_{PbTe}$) of the heterostructure, with the one on the Pb side ($\Delta Q_{Pb}$). The (4, -5) HS is highlighted in bold.